OpenStudy (anonymous):

Calculate the formula mass of each compound given below. Keep at least one decimal place in atomic masses from the periodic table.

OpenStudy (anonymous):

@zepp

OpenStudy (zepp):

?

OpenStudy (anonymous):

H2O2

OpenStudy (anonymous):

find formula mass

OpenStudy (zepp):

Peroxyde, yes

OpenStudy (anonymous):

i think i know how to do this

OpenStudy (zepp):

Mass per mole?

OpenStudy (zepp):

Simply add up the mass of each atom that peroxyde is composed of

OpenStudy (anonymous):

cuz i remember my professor did it in class

OpenStudy (anonymous):

so for H2O2 would be : 2(1.00)+2(16.0) = 2+32 = 34 amu right ?

OpenStudy (zepp):

Yep! :)

OpenStudy (anonymous):

hi @stephstar101 how r u :)

OpenStudy (stephstar101):

Hey there :) Just watchin' you learn! Mwahahahaaaa

OpenStudy (anonymous):

@zepp what about H2SO4

OpenStudy (anonymous):

lol evil laugh!!! :)))

OpenStudy (zepp):

Same thing 2(H)+(S)+4(O)

OpenStudy (anonymous):

ohh S and O is separated ?

OpenStudy (stephstar101):

Yup! They are, since they're separate elements!

OpenStudy (anonymous):

kk, this one should be 98 right?

OpenStudy (zepp):

@stephstar101 What do you think about the answer? I'm lazy to take out my calculator :<

OpenStudy (anonymous):

really now? i know you have a virtual calculator hiding somewhere !!! :O

OpenStudy (stephstar101):

You're good :) It's 34.01468 amu if you want the fancy decimal stuff

OpenStudy (zepp):

Don't want to click on icon :P

OpenStudy (stephstar101):

Oh and 98 is correct! :D

OpenStudy (anonymous):

what about "Keep at least one decimal place in atomic masses from the periodic table."

OpenStudy (anonymous):

uh oh o_0

OpenStudy (anonymous):

thanks for the freebies medal!! :D

OpenStudy (anonymous):

@zepp get your lazy self together before i smack u D:

OpenStudy (stephstar101):

I like watching people learn c: If they have the motivation to learn, they should be rewarded! :D

OpenStudy (anonymous):

your just like chatting/spamming with us...naw im just kidding lol :D

OpenStudy (zepp):

I see that @stephstar101 loves chem :D

OpenStudy (anonymous):

zepp hated chem :<

OpenStudy (anonymous):

he loves physics! BOOO !

OpenStudy (stephstar101):

:D Chem and bio are quite fun!

OpenStudy (anonymous):

just to let you guys know, i have completed 9/34 questions for a total of 4 hours! what the heck lol

OpenStudy (zepp):

I like Chem :(

OpenStudy (anonymous):

Convert 1.65 × 10^24 atoms of carbon to moles of carbon.

OpenStudy (anonymous):

uh oh here come the moles, i hate moles :<

OpenStudy (anonymous):

ohh dont u gotta use Avogadro's number for this?

OpenStudy (stephstar101):

Aww, they aren't too bad! I actually have to go soon since it's nearing midnight, but yes Avogadro all the way

OpenStudy (zepp):

http://images.wikia.com/theregularshow/images/6/6d/Mole.jpg You can't hate moles, moles are cute!

OpenStudy (anonymous):

aww wtf ! with the extra long nails lmao

OpenStudy (anonymous):

that is one ugly a$$ mole haha :))

OpenStudy (anonymous):

is this right? 2.74x10^46

OpenStudy (anonymous):

@zepp !!

OpenStudy (zepp):

?

OpenStudy (zepp):

That' wrong lol

OpenStudy (anonymous):

welcome back blitz! lol

OpenStudy (anonymous):

The number of Avogadro is used to find the number of atom in a mole \[\text{Number of atoms in a mole}=\text{Number of mole} * 6.23*10^{23}\] (I round the number)

OpenStudy (anonymous):

6.23 ? i thought its 6.02 ?

OpenStudy (anonymous):

Sorry, I left a 0 behind :s\[1.65 × 10^{24}=\text{Number of mole} * 6.023*10^{23}\] That should be our formula

OpenStudy (anonymous):

And the number of mole would be \[1.65*10^{24}\div6.023*10^{23}\]

OpenStudy (anonymous):

that's what i did =.=

OpenStudy (anonymous):

(Division, not multiplication)

OpenStudy (anonymous):

yes! i did division wth :| what did u get?

OpenStudy (zepp):

That http://puu.sh/JtvC

OpenStudy (anonymous):

i didn't put parenthesis =.=

OpenStudy (anonymous):

That must be the problem :P

OpenStudy (anonymous):

How many moles of Fe2O3 are in 235 g of the compound?

OpenStudy (anonymous):

ok let me try this first

OpenStudy (anonymous):

Cool :)

OpenStudy (anonymous):

idk if i did it right or not, but my final answer is 1.47

OpenStudy (anonymous):

Alright so in Fe2O3, there's two Fe and 3 O 2(Fe)= 55.845*2 = 111.69 3(O)= 16*3 = 48 Fe2O3 = 111.69 + 48 = 159.69 Take the 235 and divide 235/159.69 = 2.1

OpenStudy (anonymous):

1.47, sorry, not 2.1

OpenStudy (anonymous):

So yep, that's right :)

OpenStudy (anonymous):

:) thanks for the explanation !

OpenStudy (zepp):

yw :)

OpenStudy (anonymous):

not you, i was thanking Blitz ! lol

OpenStudy (zepp):

psh

OpenStudy (anonymous):

You are welcome :D

OpenStudy (anonymous):

Blitz is a really helpful guy, he knows everything well, explain clearly, give good examples..everything you can ask of a tutors :0

OpenStudy (anonymous):

blitz is awesome :)

OpenStudy (zepp):

So I don't explain well? :(

OpenStudy (anonymous):

zepp you just lazyyy !! 50% of blitz muahahha :))

OpenStudy (zepp):

meh

OpenStudy (anonymous):

@zepp I thought you need to be helpful here and not just talking rubbish... :/

OpenStudy (anonymous):

yup ! he need to be more helpful and not disconnecting too much! and also get a new ISP lol

OpenStudy (anonymous):

Next question? :D

OpenStudy (anonymous):

lol :) your so hyper

OpenStudy (anonymous):

How many moles of N are in 0.235 g of N2O?

OpenStudy (anonymous):

i got this, just need you to chek my answer ok?

OpenStudy (zepp):

42 is the answer.

OpenStudy (anonymous):

wow...i said hold on lmao ! -.-

OpenStudy (anonymous):

wtf zepp ! screww u lol :D

OpenStudy (zepp):

:D

OpenStudy (anonymous):

i got .0083 though ?

OpenStudy (anonymous):

It's NO2 or N2O?

OpenStudy (anonymous):

N2O !

OpenStudy (anonymous):

Alright, first step would be see how many g does 1 mole of this weight 2(N) = 14 1(O) = 16 N2O = 14+16 = 30 30g/mol of N2O We have 0.235 g We'll simply divide, how many 30g/mol fits in there? \(0.235g\div30g/mol=0.007833mol\) You forgot a 7 somewhere

OpenStudy (anonymous):

wait does the O even necessarily in this?

OpenStudy (anonymous):

it only ask for "N"

OpenStudy (zepp):

omg I misread the question, hold on

OpenStudy (anonymous):

.....

OpenStudy (zepp):

What @Blitzkrieg did was correct, there's 2 moles of N in 30g.

OpenStudy (zepp):

Since there's 0.007833mol of N2O that fits in 0.235 and there's 2 moles of N in 1 mole of N2O, you multiply that number by 2

OpenStudy (anonymous):

2x14=28

OpenStudy (anonymous):

and take .235/28 = .0084